# Continuum Attention for Neural Operators

### v3

@@ 3.2 Patched Cross-Attention
[…]
@@ 4 Transformer Neural Operators
In this section we turn our focus to describing how to devise transformer-based neural operators using the continuum attention operator $\mathsf{A}$ as defined in the function space setting in Section 2, and the continuum patched-attention operator $\mathsf{A}_{\mathsf{P}}$ acting on function space from Section 3. We make use of the operator frameworks developed in Sections 2 and 3 to formulate transformer neural operator architectures acting on function space. We start in Subsection 4.1, setting-up the framework. In Subsection 4.2 we formulate an analogue of the transformer from Vaswani et al. (2017) as an operator mapping an input function space $\mathcal{U}\bigl{(}D;\mathbbm{R}^{d_{u}}\bigr{)}$ to the solution function space $\mathcal{Z}\bigl{(}D;\mathbbm{R}^{d_{z}}\bigr{)}$. The resulting scheme, which we define as transformer neural operator (TNO) is invariant to the discretization of the input function and allows for zero-shot generalization to irregular, non-uniform meshes. However, this architecture exhibits quadratic complexity with respect to the number of discretization points of the input functions. Hence, motivated by the need to develop new efficient architectures to employ attention on longer sequences, or in our case higher resolution discretizations of functions, in Subsection 4.3 we describe a generalization of the ViT methodology to the function space setting, inspired by the work in Dosovitskiy et al. (2021). In particular we generalize the procedure of lifting flattened patches to the embedding dimension, as done in ViT, to the function space setting. This is achieved by lifting patches of functions to an embedding dimension using a nonlocal linear operator, namely, an integral operator. Unlike ViT, the resulting neural operator architecture, which we call ViT neural operator (ViTNO), is discretization invariant and allows for zero-shot generalization to different resolutions. In Subsection 4.4 we describe a different patching-based transformer neural operator acting on function space, the Fourier attention neural operator (FANO). The architecture is discretization invariant and is based on $\mathsf{Q},\mathsf{K},\mathsf{V}$, learnable parameters in the attention mechanism, being defined as integral operators acting themselves on function space. We demonstrate in Section 6 that the approach involving parametrizing $\mathsf{Q},\mathsf{K},\mathsf{V}$ as integral operators yields an architecture with greater parameter count for a similar computational complexity. We show that this architecture produces improved results in the context of smooth problems.
@@ 4.1 Set-Up
for any $u\in\mathcal{U}(D;\mathbbm{R}^{d_{u}})$ and $\theta\in\Theta$. In (9), the operators $\mathsf{T}_{\textrm{in}}:\mathcal{U}(D;\mathbbm{R}^{d_{u}})\times\Theta\to\mathcal{V}$ and $\mathsf{T}_{\textrm{out}}:\mathcal{V}\times\Theta\to\mathcal{Z}(D;\mathbbm{R}^{d_{z}})$ are defined for some appropriate embedding function space $\mathcal{V}$. We will make explicit the definitions of the general operators $\mathsf{T}_{\textrm{in}}$, $\mathsf{T}_{\textrm{out}}$ and the embedding function space $\mathcal{V}$ in the context of each architecture. Throughout this section, we drop explicit dependencies on $\theta$ for brevity unless we wish to stress a parametric dependence; for example, we may abuse notation and write $\mathsf{T}_{\textrm{in}}(u):=\mathsf{T}_{\textrm{in}}(u;\ \theta)$. The operator $\mathsf{E}_{L}:\mathcal{V}\times\Theta\to\mathcal{V}$ defines the neural operator analogue of the transformer encoder block from Vaswani et al. (2017). Its action on the input $v^{(0)}\coloneqq\mathsf{T}_{\textrm{in}}(u)\in\mathcal{V}$ is summarized by the iteration
@@
Assuming ${\kappa}_{\theta}(x,y)={\kappa}_{\theta}(x-y)$, hence the kernel to be periodic, then (12) may be viewed as a convolution operator, which makes it possible to use the Fast Fourier Transform (FFT) to compute (12) and to parametrize $\mathsf{W}$ in Fourier space. Following Li et al. (2021), this insight leads to the following formulation of the operator given in Definition 19.
@@ 4.3 Vision Transformer Neural Operator
We introduce an analogue of the vision transformer architecture as a neural operator using the patching in the continuum framework developed in Section 3. We note that in the context of the ViT from Dosovitskiy et al. (2021), performing patch “flattening” before a local lifting to the embedding dimension introduces a nonlocal transformation on the patches, as described in Remark 15. This procedure as implemented in Dosovitskiy et al. (2021) violates the desirable (for neural operators) discretization invariance property. Hence, for a neural operator analogue, we substitute the lifting to embedding dimension with a nonlocal integral operator. The ViT neural operator (ViT NO) introduced is thus discretization invariant. We describe this methodology within the framework provided by making explicit choices for $\mathsf{T}_{\textrm{in}},\mathsf{T}_{\textrm{out}},\mathsf{E}_{L}$ that instantiate Equations 9 and 10 appropriately. In Figure 2 we provide a schematic for the architecture, with a graphical representation of the encoder layer in Figure 3.
@@ 4.4 Fourier Attention Neural Operator
[…]
@@ 5 Universal Approximation by Transformer Neural Operators
Universal approximation is a minimal requirement for neural operators. When the operator to be approximated maps between spaces of functions defined over Euclidean domains, universal approximation necessarily requires both nonlocality and nonlinearity. In the recent paper Lanthaler et al. (2025), a minimal architecture possessing these two properties is exhibited. In this section, rather than trying to prove universal approximation for the various discretization-invariant neural operators introduced in the paper, we construct a simple canonical setting that exhibits the universality of the attention mechanism on function space, allowing direct exploitation of the ideas in Lanthaler et al. (2025). We build on the notation used in Subsection 4.1. In particular $\mathsf{A}$ denotes the self-attention operator from Definition 5 and $f$ an activation function. Throughout this section, we consider activation functions $f\in C^{\infty}(\mathbbm{R})$ which are non-polynomial and Lipschitz continuous.
@@
for any $x\in D$. Note that the neural operator thus defined does not include layer normalizations; it is thus a variant of the transformer neural operator from Subsection 4.2. We may now apply the result of Lanthaler et al. (2025) to show two universal approximation theorems for the resulting transformer neural operator.
@@
The existence of $W_{1},R_{1},R_{2},P_{1},P_{2},b_{1},b_{R},b_{R}^{\prime},b_{P},b_{P}^{\prime}$ so that $\mathsf{G}(\cdot\,;\theta)$ satisfies (35) then follows from Lanthaler et al. (2025, Theorem 2.1), which also involves the application of the universality result for two-layer neural networks of Pinkus (1999, Theorem 4.1).
The analysis in Lanthaler et al. (2025) allows the derivation of an analogous universal approximation theorem for functions belonging to Sobolev spaces. This generalization is the content of the next theorem. The proof follows easily by applying the same argument as in the proof of Theorem 22 and the result of Lanthaler et al. (2025, Theorem 2.2).
@@ 6 Numerical Experiments
In this subsection we consider problems where attention is applied to functions in the one-dimensional time domain. The first setting we consider is that of the Lorenz ‘63 dynamical system (Lorenz, 1963), a simplified model for atmospheric convection. The dynamics are given by
@@
This operator would only be well-defined if $(y(0),z(0))$ were included as inputs. We nonetheless expect accurate recovery of the hidden trajectory by the property of synchronization (Pecora and Carroll, 1990; Hayden et al., 2011).
@@
The experimental settings we consider for the continuous time operator learning problem may be viewed through the lens of data assimilation as smoothing problems (Sanz-Alonso et al., 2023). Indeed, the mappings to be learned represent the recovery of an unobserved trajectory $\{z(t)\}_{t\in{[0,T]}}$ conditioned on the observation of a whole trajectory $\{u(t)\}_{t\in{[0,T]}}$.
@@
In other words, the second half of the testing domain is up-sampled by a factor of 2 compared to the training discretization. As seen qualitatively in Figure 8 testing on the non-uniform discretization based on a model trained on the uniform grid is successful. The use of different grids in test and train does lead to a larger error than the one incurred when they match (Figure 7), but the errors remain small. It is our continuum formulation of attention from Section B.1 that enables this direct generalization; indeed, as displayed in Figure 9, the transformer from Vaswani et al. (2017) when applied to the same setting yields larger errors than the transformer neural operator, due to the TNO’s reweighting based on the discretization of the grid.
@@
Figure 9: The panel displays the performance in relative $L^{2}$ errors of the transformer from Vaswani et al. (2017) when applied to the Lorenz ‘63 operator learning problem of recovering the unobserved $y$ trajectory. The results displayed concern a model trained on trajectories indexed on the index set (42) deployed on a test of trajectories indexed on (41). In each column, the top plot shows the input data and the bottom compares the prediction and ground truth; the left column shows the testing example achieving a median error, and the right columns shows the test example achieving the worst error.
@@ 6.2.1 Results
The results presented concern the two variants of Darcy flow, with lognormal and piecewise constant inputs, and the Kolmogorov flow. The operator learning architectures we deploy include the three transformer-based methods introduced in this paper, and the implementations of the Fourier neural operator from the “NeuralOperator” library (Li et al., 2021; Kovachki et al., 2023), the Galerkin transformer from the “galerkin-transformer” library (Cao, 2021), and the AFNO (Guibas et al., 2022) from the “physics-nemo” library (PhysicsNeMo Contributors, 2023). We note that in order to obtain a numerical comparison of the attention mechanisms themselves, in the context of the Galerkin transformer architecture we only use the Galerkin transformer encoder, with lifting and projections achieved by linear layers as in TNO. This allows for a more direct comparison between the Fourier attention from Cao (2021) and our continuum attention, without potential expressivity gain from additional architectural components. In the following discussion, we do not include experiments performed using the Codomain-attention neural operator (Rahman et al., 2024) in the high-resolution settings: the memory usage with the implementation from the “NeuralOperator” library did not allow for comparably sized models, in terms of parameters; the performance of the models we trained was hence incomparably inferior to the others.
@@
We now consider training two patched architectures on the Darcy Flow problem with piecewise constant inputs and the Kolmogorov Flow problem; in both cases we use a $416\times 416$ resolution setting. We compare cost versus accuracy for the different neural operators implemented, defining cost in terms of both number of parameters and number of FLOPS. In Table 2 we present the FLOPS for each method; details of the derivation are provided in Appendix C. We note that the parameter scaling for the transformer neural operator is given by $\mathcal{O}\bigl{(}d_{\textrm{model}}^{2}\bigr{)}$, while the scaling for the FNO and the Fourier attention neural operator is $\mathcal{O}\bigl{(}d_{\textrm{model}}^{2}\cdot k_{\textrm{max}}\bigr{)}$, where $k_{\textrm{max}}$ is the total number of Fourier modes used. The parameter scaling for the ViT neural operator is given by $\mathcal{O}\bigl{(}d_{\textrm{model}}\cdot k_{\textrm{max}}+d_{\textrm{model}}^{2}\bigr{)}$. We note that the parameter scaling for the AFNO architecture is given by $\mathcal{O}\bigl{(}Nd_{\textrm{model}}+d_{\textrm{model}}^{2}\bigr{)}$, where $b$ is a chosen integer hyperparameter, and where the $Nd_{\textrm{model}}$ appears because of the learnable positional encoding. Further details are provided in Appendix C.
@@
Table 2: Evaluation cost (measured in FLOPS) for the three proposed transformer neural operators compared to the Fourier neural operator (Li et al., 2021) and Adaptive Fourier neural operator (Guibas et al., 2022). We include the scaling of the AFNO as implemented for our use case, i.e. not employing patching and not employing mode truncation. We give further details on the derivation of these scalings in Appendix C.
Figures 10 and 11 display the cost-accuracy trade-off. Table 3 provides the test relative $L^{2}$ errors values, depicted in the cost-accuracy analysis figures, obtained with the models with the highest parameter count. Figure 10 demonstrates the parameter-efficiency of the ViT neural operator in comparison with the FNO, AFNO and Fourier attention neural operator. Once cost is defined in terms of FLOPS it is apparent that both patched transformer architectures outperform the FNO; furthermore in the Kolmogorov flow problem, the FNO exhibits accuracy-saturation, caused by the finite data set, whereas the patched architectures are less affected, in the ranges we test here. This suggests that the patched transformer architectures use the information content in the data more efficiently. The AFNO is comparable in terms of parameter count to FANO but more costly than ViTNO. On the other hand, even with no patching, the AFNO is cheaper in terms of evaluation complexity. Both ViTNO and FANO outperform the AFNO in terms of parameter-accuracy tradeoff in the context of the Darcy flow problem. For the same problem ViTNO and FANO bring an improvement in accuracy, albeit at a higher computational complexity. In the context of the Kolmogorov flow problem, ViTNO and FANO are competitive with AFNO in parameter-accuracy tradeoff. However for this problem, the AFNO can achieve higher accuracy with a lower complexity. It is important to note that because of the learnable positional embedding, and potential strided convolutions (if using patching), the AFNO is not a neural operator. This means that applying this architecture to different resolutions would require retraining, thus significantly increasing computational complexity. This is a fundamental advantage of the other architectures as their parameters may be deployed across different resolutions for training and inference. We recall the example of GenCast (Price et al., 2025), where a medium-range weather forecasting model is trained using a neural operator at low resolutions, and then finetuned at higher resolutions, thus drastically reducing the cost for such a large scale application.

### v4

@@ 3.2 Patched Cross-Attention
$\widetilde{v}(o)={\left.\kern-1.2ptv\mathchoice{\vphantom{\big|}}{}{}{}\right|_{E_{o}}},$ (8)
@@ 4 Transformer Neural Operators
In this section we turn our focus to describing how to devise transformer-based neural operators using the continuum attention operator $\mathsf{A}$ as defined in the function space setting in Section 2, and the continuum patched-attention operator $\mathsf{A}_{\mathsf{P}}$ acting on function space from Section 3. We make use of the operator frameworks developed in Sections 2 and 3 to formulate transformer neural operator architectures acting on function space. We start in Subsection 4.1, setting-up the framework. In Subsection 4.2 we formulate an analogue of the transformer from vaswani2017attention as an operator mapping an input function space $\mathcal{U}\bigl(D;\mathbbm{R}^{d_{u}}\bigr)$ to the solution function space $\mathcal{Z}\bigl(D;\mathbbm{R}^{d_{z}}\bigr)$. The resulting scheme, which we define as transformer neural operator (TNO) is invariant to the discretization of the input function and allows for zero-shot generalization to irregular, non-uniform meshes. However, this architecture exhibits quadratic complexity with respect to the number of discretization points of the input functions. Hence, motivated by the need to develop new efficient architectures to employ attention on longer sequences, or in our case higher resolution discretizations of functions, in Subsection 4.3 we describe a generalization of the ViT methodology to the function space setting, inspired by the work in dosovitskyi2021image. In particular we generalize the procedure of lifting flattened patches to the embedding dimension, as done in ViT, to the function space setting. This is achieved by lifting patches of functions to an embedding dimension using a nonlocal linear operator, namely, an integral operator. Unlike ViT, the resulting neural operator architecture, which we call ViT neural operator (ViTNO), is discretization invariant and allows for zero-shot generalization to different resolutions. In Subsection 4.4 we describe a different patching-based transformer neural operator acting on function space, the Fourier attention neural operator (FANO). The architecture is discretization invariant and is based on $\mathsf{Q},\mathsf{K},\mathsf{V}$, learnable parameters in the attention mechanism, being defined as integral operators acting themselves on function space. We demonstrate in Section 6 that the approach involving parametrizing $\mathsf{Q},\mathsf{K},\mathsf{V}$ as integral operators yields an architecture with greater parameter count for a similar computational complexity. We show that this architecture produces improved results in the context of a problem with smooth inputs.
@@ 4.1 Set-Up
for any $u\in\mathcal{U}(D;\mathbbm{R}^{d_{u}})$ and $\theta\in\Theta$. In (9), the operators $\mathsf{T}_{\textrm{in}}:\mathcal{U}(D;\mathbbm{R}^{d_{u}})\times\Theta\to\mathcal{V}$ and $\mathsf{T}_{\textrm{out}}:\mathcal{V}\times\Theta\to\mathcal{Z}(D;\mathbbm{R}^{d_{z}})$ are defined for some appropriate embedding function space $\mathcal{V}$. We will make explicit the definitions of the general operators $\mathsf{T}_{\textrm{in}}$, $\mathsf{T}_{\textrm{out}}$ and the embedding function space $\mathcal{V}$ in the context of each architecture. Throughout this section, we drop explicit dependencies on $\theta$ for brevity unless we wish to stress a parametric dependence; for example, we may abuse notation and write $\mathsf{T}_{\textrm{in}}(u):=\mathsf{T}_{\textrm{in}}(u;\ \theta)$. The operator $\mathsf{E}_{L}:\mathcal{V}\times\Theta\to\mathcal{V}$ defines the neural operator analogue of the transformer encoder block from vaswani2017attention. Its action on the input $v^{(0)}\coloneqq\mathsf{T}_{\textrm{in}}(u)\in\mathcal{V}$ is summarized by the iteration
@@
Assuming ${\kappa}_{\theta}(x,y)={\kappa}_{\theta}(x-y)$, hence the kernel to be periodic, then (12) may be viewed as a convolution operator, which makes it possible to use the Fast Fourier Transform (FFT) to compute (12) and to parametrize $\mathsf{W}$ in Fourier space. Following li2021fourier, this insight leads to the following formulation of the operator given in Definition 19.
@@ 4.3 Vision Transformer Neural Operator
We introduce an analogue of the vision transformer architecture as a neural operator using the patching in the continuum framework developed in Section 3. We note that in the context of the ViT from dosovitskyi2021image, performing patch “flattening” before a local lifting to the embedding dimension introduces a nonlocal transformation on the patches, as described in Remark 15. This procedure as implemented in dosovitskyi2021image violates the desirable (for neural operators) discretization invariance property. Hence, for a neural operator analogue, we substitute the lifting to embedding dimension with a nonlocal integral operator. The ViT neural operator (ViT NO) introduced is thus discretization invariant. We describe this methodology within the framework provided by making explicit choices for $\mathsf{T}_{\textrm{in}},\mathsf{T}_{\textrm{out}},\mathsf{E}_{L}$ that instantiate Equations 9 and 10 appropriately. In Figure 2 we provide a schematic for the architecture, with a graphical representation of the encoder layer in Figure 3.
@@ 4.4 Fourier Attention Neural Operator
for any $j\in\llbracket 1,P\rrbracket$ and any $x\in D$, where ${u}_{\textrm{in}}\in\mathcal{U}(D;\mathbbm{R}^{d_{u}+d})$ is defined as in (14), $\mathsf{S}_{\textrm{reshape}}$ as in (20) and the nonlocal operation (21) is replaced by a local pointwise one.
@@ 5 Universal Approximation by Transformer Neural Operators
Universal approximation is a minimal requirement for neural operators. When the operator to be approximated maps between spaces of functions defined over Euclidean domains, universal approximation necessarily requires both nonlocality and nonlinearity. In the recent paper lanthaler2023nonlocal, a minimal architecture possessing these two properties is exhibited. In this section, rather than trying to prove universal approximation for the various discretization-invariant neural operators introduced in the paper, we construct a simple canonical setting that exhibits the universality of the attention mechanism on function space, allowing direct exploitation of the ideas in lanthaler2023nonlocal. We build on the notation used in Subsection 4.1. In particular $\mathsf{A}$ denotes the self-attention operator from Definition 5 and $f$ an activation function. Throughout this section, we consider activation functions $f\in C^{\infty}(\mathbbm{R})$ which are non-polynomial and Lipschitz continuous.
@@
for any $x\in D$. Note that the neural operator thus defined does not include layer normalizations; it is thus a variant of the transformer neural operator from Subsection 4.2. We may now apply the result of lanthaler2023nonlocal to show two universal approximation theorems for the resulting transformer neural operator.
@@
The existence of $W_{1},R_{1},R_{2},P_{1},P_{2},b_{1},b_{R},b_{R}^{\prime},b_{P},b_{P}^{\prime}$ so that $\mathsf{G}(\cdot\,;\theta)$ satisfies (35) then follows from lanthaler2023nonlocal, which also involves the application of the universality result for two-layer neural networks of Pinkus_1999.
The analysis in lanthaler2023nonlocal allows the derivation of an analogous universal approximation theorem for functions belonging to Sobolev spaces. This generalization is the content of the next theorem. The proof follows easily by applying the same argument as in the proof of Theorem 22 and the result of lanthaler2023nonlocal.
@@ 6 Numerical Experiments
In this subsection we consider problems where attention is applied to functions in the one-dimensional time domain. The first setting we consider is that of the Lorenz ‘63 dynamical system (lorenz1963deterministic), a simplified model for atmospheric convection. The dynamics are given by
@@
This operator would only be well-defined if $(y(0),z(0))$ were included as inputs. We nonetheless expect accurate recovery of the hidden trajectory by the property of synchronization (pecora1990synchronization; hayden2011discrete).
@@
The experimental settings we consider for the continuous time operator learning problem may be viewed through the lens of data assimilation as smoothing problems (Sanz-Alonso_Stuart_Taeb_2023). Indeed, the mappings to be learned represent the recovery of an unobserved trajectory $\{z(t)\}_{t\in{[0,T]}}$ conditioned on the observation of a whole trajectory $\{u(t)\}_{t\in{[0,T]}}$.
@@
In other words, the second half of the testing domain is up-sampled by a factor of 2 compared to the training discretization. As seen qualitatively in Figure 8 testing on the non-uniform discretization based on a model trained on the uniform grid is successful. The use of different grids in test and train data does lead to a larger error than the one incurred when they match (Figure 7), but the errors remain small. It is our continuum formulation of attention from Section B.1 that enables this direct generalization; indeed, as displayed in Figure 9, the transformer from vaswani2017attention when applied to the same setting yields larger errors than the transformer neural operator, due to the TNO’s reweighting based on the discretization of the grid.
@@
Figure 9: The panel displays the performance in relative $L^{2}$ errors of the transformer from vaswani2017attention when applied to the Lorenz ‘63 operator learning problem of recovering the unobserved $y$ trajectory. The results displayed concern a model trained on trajectories indexed on the index set (42) deployed on a test of trajectories indexed on (41). In each column, the top plot shows the input data and the bottom compares the prediction and ground truth; the left column shows the testing example achieving a median error, and the right columns shows the test example achieving the worst error.
@@ 6.2.1 Results
The results presented concern the two variants of Darcy flow, with lognormal and piecewise constant inputs, and the Kolmogorov flow. The operator learning architectures we deploy include the three transformer-based methods introduced in this paper, and the implementations of the Fourier neural operator from the “NeuralOperator” library (li2021fourier; kovachki2023neural), the Galerkin transformer from the “galerkin-transformer” library (cao2021choose), and the AFNO (guibas2022efficient) from the “physics-nemo” library (physicsnemo2023). We note that in order to obtain a numerical comparison of the attention mechanisms themselves, in the context of the Galerkin transformer architecture we only use the Galerkin transformer encoder, with lifting and projections achieved by linear layers as in TNO. This allows for a more direct comparison between the Fourier attention from cao2021choose and our continuum attention, without potential expressivity gain from additional architectural components. In the following discussion, we do not include experiments performed using the Codomain-attention neural operator (rahman2024pretraining) in the high-resolution settings: the memory usage with the implementation from the “NeuralOperator” library did not allow for comparably sized models, in terms of parameters; the performance of the models we trained was hence incomparably inferior to the others.
@@
We now consider training two patched architectures on the Darcy Flow problem with piecewise constant inputs and the Kolmogorov Flow problem; in both cases we use a $416\times 416$ resolution setting. We compare cost versus accuracy for the different neural operators implemented, defining cost in terms of both number of parameters and number of FLOPS. In Table 2 we present the FLOPS for each method; details of the derivation are provided in Appendix C. We note that the parameter scaling for the transformer neural operator is given by $\mathcal{O}\bigl(d_{\textrm{model}}^{2}\bigr)$, while the scaling for the FNO and the Fourier attention neural operator is $\mathcal{O}\bigl(d_{\textrm{model}}^{2}\cdot k_{\textrm{max}}\bigr)$, where $k_{\textrm{max}}$ is the total number of Fourier modes used. The parameter scaling for the ViT neural operator is given by $\mathcal{O}\bigl(d_{\textrm{model}}\cdot k_{\textrm{max}}+d_{\textrm{model}}^{2}\bigr)$. We note that the parameter scaling for the AFNO architecture is given by $\mathcal{O}\bigl(Nd_{\textrm{model}}+d_{\textrm{model}}^{2}\bigr)$, where the $Nd_{\textrm{model}}$ appears because of the learnable positional encoding. Further details are provided in Appendix C.
@@
Table 2: Evaluation cost (measured in FLOPS) for the three proposed transformer neural operators compared to the Fourier neural operator (li2021fourier) and adaptive Fourier neural operator (guibas2022efficient). We include the scaling of the AFNO as implemented for our use case, i.e. not employing patching and not employing mode truncation; we further note that in the context of this architecture $b$ is a chosen integer hyperparameter. We give further details on the derivation of these scalings in Appendix C.
Figures 10 and 11 display the cost-accuracy trade-off. Table 3 provides the test relative $L^{2}$ errors values, depicted in the cost-accuracy analysis figures, obtained with the models with the highest parameter count. Figure 10 demonstrates the parameter-efficiency of the ViT neural operator in comparison with the FNO, AFNO and Fourier attention neural operator. Once cost is defined in terms of FLOPS it is apparent that both patched transformer architectures outperform the FNO; furthermore in the Kolmogorov flow problem, the FNO exhibits accuracy-saturation, caused by the finite data set, whereas the patched architectures are less affected, in the ranges we test here. This suggests that the patched transformer architectures use the information content in the data more efficiently. The AFNO is comparable in terms of parameter count to FANO but more costly than ViTNO. On the other hand, even with no patching, the AFNO is cheaper in terms of evaluation complexity. Both ViTNO and FANO outperform the AFNO in terms of parameter-accuracy tradeoff in the context of the Darcy flow problem. For the same problem ViTNO and FANO bring an improvement in accuracy, albeit at a higher computational complexity. In the context of the Kolmogorov flow problem, ViTNO and FANO are competitive with AFNO in parameter-accuracy tradeoff. However for this problem, the AFNO can achieve higher accuracy with a lower complexity. It is important to note that because of the learnable positional embedding, and potential strided convolutions (if using patching), the AFNO is not a neural operator. This means that applying this architecture to different resolutions would require retraining, thus significantly increasing computational complexity. This is a fundamental advantage of the other architectures as their parameters may be deployed across different resolutions for training and inference. We recall the example of GenCast (Price2025), where a medium-range weather forecasting model is trained using a neural operator at low resolutions, and then finetuned at higher resolutions, thus drastically reducing the cost for such a large scale application.